# From Planetesimals to Dwarf Planets by Pebble Accretion

### v1

@@ 4.1 Formation in a pebble ring
Pebble rings are ubiquitous in ALMA imagery (Andrews et al., 2018; Bae et al., 2023). Containment of particles by pressure maxima is the leading explanation for the observed particles accumulation in these rings. Around the pressure maximum location $\eta\approx 0$ and the relative motion between pebbles and bigger bodies is due to turbulence. Studies have used the extent of rings in both the vertical and radial dimensions to constrain the diffusivity—the “turbulent alpha”—of the gas. Specifically, this constrains the ratio of gas diffusivity over Stokes number of the particles, $\alpha_{D}/\tau_{s}$. Rosotti (2023) compiled results from a number of works that measure the radial width of continuum rings (their Table 2), finding that the inferred $\alpha_{D}/\tau_{s}$ ranged from a low of 0.04 to values exceeding unity (the latter essentially implies particles follow the pressure distribution). However, $\alpha_{D}/\tau_{s}$ obtained from the vertical extent of the dust layer are often inferred to be smaller, e.g., $\alpha_{D}/\tau_{s}\approx 10^{-2}$ (Villenave et al., 2022). For simplicity, we adopt in the following a range in $\alpha_{D}/\tau_{s}$ between $0.01$ and $0.4$, which covers most of these measurements.
When we convert the diffusivity parameter $\alpha_{D}$ to a root-mean-square (rms) particle velocity, we obtain [^10] Here we have related the parameter $\alpha_{D}$, which expresses the diffusivity of the gas, to the rms velocity of the turbulent gas motions $\sigma$. The turbulent diffusivity of the gas is parameterized in terms of $\alpha$ as in $D_{\mathrm{gas}}=\alpha_{D}c_{s}^{2}/\Omega$. In general, we can write $D_{\mathrm{gas}}=V_{L}^{2}t_{L}$ where $V_{L}$ and $t_{L}$ are, respectively, the velocities and turnover times of the largest eddies in the turbulent cascade. A common choice for rotating systems is $t_{L}=\Omega_{K}^{-1}$ and therefore $V_{L}=\sqrt{\alpha_{D}}c_{s}$ (Cuzzi et al., 2001). Furthermore, if turbulence is Kolmogorov, the rms velocities of the turbulent gas is $\Delta v=\mathcal{M}c_{s}=(3/2)V_{L}$. Then, we obtain $\alpha_{D}=(2\mathcal{M}/3)^{2}=(2\sigma/3h)^{2}$.
@@
Figure 8: Constraints on the particle aerodynamic size (Stokes number $\tau_{s}$) and turbulence intensity of the gas, where $\sigma$ is identified as the root-mean-square turbulent gas velocity. Constraints follow from simulating the TNO size distribution (run turb-1p, black line), ALMA ring morphological analysis (blue lines; Rosotti 2023), and particle fragmentation velocity (orange lines; the adopted value for the Keplerian velocity $v_{K}$ is the circular motion at 20 au). Under the assumption that the primordial belt featured conditions similar to ALMA rings the degeneracy between $\tau_{s}$ and $\sigma$ (or $\alpha_{D}$; Equation (16)) can be lifted and we find, $\tau_{s}\sim 0.01$, $\alpha_{D}\sim 10^{-3}$, and a fragmentation threshold of ${\sim}2\,\mathrm{m\,s}^{-1}$.
A further constraint to particle sizes in rings comes from the fragmentation barrier (Güttler et al., 2010; Birnstiel et al., 2011). When the gas is turbulent, the pebble-pebble relative velocity increases with the size (Stokes number) as in $\Delta v_{\mathrm{pp}}\approx\mathcal{M}c_{s}\sqrt{2\tau_{s}}$ (Ormel & Cuzzi, 2007), [^11] This is the rms particle-particle relative velocity between two particles in the inertial range. The rms particle-planetesimal relative velocity for $\tau_{s}<1$ particles is $\Delta v=\mathcal{M}c_{s}$. where $\mathcal{M}$ is the Mach number of the turbulent flow. The aerodynamic size is then determined by the condition $\Delta v_{\mathrm{pp}}=v_{\mathrm{frag}}$, resulting in:
@@
These constraints are presented in Figure 8, together with the constraint on $\tau_{s}\sigma^{3}$ that we obtained from modelling the TNO size distribution (Table 2). The constraint on the extent of ALMA rings are almost orthogonal to the constraint from the TNO size distribution modelling (this study), allowing us to put meaningful constraints on $\tau_{s}$ and $\alpha_{D}$. Together, they point to an aerodynamic size in the range of $\tau_{s}\sim[3\times 10^{-3}{-}4\times 10^{-2}]$, a turbulent diffusivity of $\alpha_{D}\sim[4\times 10^{-4}{-}2\times 10^{-3}]$, and a fragmentation threshold of $v_{\mathrm{frag}}\sim[1.7{-}3.2\,\mathrm{m\,s}^{-1}]$.
These values are similar to what was found by Jiang et al. (2024) in a recent analysis, who assumed a fragmentation threshold of $v_{\mathrm{frag}}\sim 1\,\mathrm{m\,s}^{-1}$. Our results independently suggest that the particle fragmentation threshold amount is ${\sim}2\,\mathrm{m\,s}^{-1}$ in order for particles to grow to the inferred aerodynamic sizes. (A higher fragmentation threshold is possible if the growth of particles is limited by, e.g., bouncing; Zsom et al. 2010.) A fragmentation velocity increase by a factor of 2 would increase the derived value of the fragmentation $\alpha$ in Jiang et al. (2024) by a factor of 4 to ${\sim}4\times 10^{-4}$, in line with our findings.
@@ 4.2 A laminar disk
If pebbles are not entrained by a pressure maximum and freely drift inward, the relative velocity between planetesimals and pebbles is given by the global disk pressure gradient $\eta$. From Equation (8), $\sigma=\eta\approx 5\times 10^{-3}$ for a power-law index of pressure on radius of $2$ and from the $\tau_{s}$$\sigma$ relationship (Equation (15)), we then find an aerodynamic size of a mere $\tau_{s}=7\times 10^{-4}$, which is rather small for pebbles. Indeed, the radial drift timescale
@@
already rivals the disk lifetime, suggesting that particles hardly drift. In addition, growth of planetesimals by such small particles would be excruciatingly slow, $t_{\mathrm{gr}}\sim t_{0}/\tau_{s}\sim 50\,\mathrm{Myr}$ for the parameters listed in Equation (5), far exceeding the lifetime of the gas disk. A growth timescale ${\lesssim}10^{6}$ yr would therefore imply a (midplane) dust-to-gas ratio $Z_{\mathrm{peb}}\sim 1$, which seems unlikely for small $\tau_{s}$ particles. In summary, it is unlikely that TNOs could accrete pebbles from an inward-drifting pebble stream in the standard, laminar disk model.
In addition, pebble accretion in laminar disk is lossy (Ormel, 2017; Lin et al., 2018). In the scenario where the birth ring is supported by pressure, all pebbles can (eventually) be accreted. That is, the efficiency is near 100%. This contrasts with the smooth disk scenario, where pebbles that fail to be accreted are lost. Consequently, the total mass in pebbles required for accretion, $m_{\mathrm{needed}}$, exceeds the mass in pebbles actually accreted ($m_{\mathrm{peb}}$). In a laminar disk, we can estimated $m_{\mathrm{needed}}$ as
@@
All these concerns would alleviate considerably when the disk is colder. Expressed in terms of the disk aspect ratio ($h\propto T^{1/2}$), $\eta\propto h^{2}$ and therefore $\tau_{s}\propto\sigma^{-3}\propto h^{-6}$ from Equation (15). For example, when $h=0.05$ (about 35 K at 20 au; see Equation (7)), $\tau_{s}\approx 6\times 10^{-3}$, a factor of 10 increase, the drift timescale reduces by a factor of 4 compared to Equation (19), the growth timescale may fall below the lifetime of the disk, and the require pebble mass $m_{\mathrm{needed}}$ to values below $500\,m_{\oplus}$. These numbers are within the realm of pebble accretion models (Lambrechts et al., 2019) albeit on the high side. A question that arises, however, is where these $500\,m_{\oplus}$ in solids would end up, as the Solar System’s giant planets, while enriched, are not dominated in metals.
@@ 5.2 Formation time constraints
Recently, Cañas et al. (2024) also investigated the growth of TNOs by pebble accretion. Their approach is complementary to ours. In particular, their study emphasizes the composition and internal density of TNOs. A key constraint is that ”small” TNOs of radius $50{-}100$ km could not have formed too quickly as radiogenic heating—primarily Al-26 decay at a half-life of 0.7 Myr (Norris et al., 1983)—would have molten these bodies and reduced their porosity. This outcome contrasts observations, which indicates that these TNOs have retained significant porosity (Bierson & Nimmo, 2019).
@@
Cañas et al. (2024) addressed this conundrum by hypothesizing that the initial distribution of TNOs formed through streaming instabilities dominated by large, ice-rich pebbles. Processes as photodesorption could have resulted in a compositional dichotomy where ice is effectively transferred to the larger particles, which settled to the midplane layers that were shielded from stellar UV irradiation. Cañas et al. (2024) argue that streaming instability would favor the larger, ice-rich pebbles over the smaller Al-26 rich ones. Hence, the first generation of planetesimals were low in Al-26 and the early heating problem outlined by Bierson & Nimmo (2019) is avoided. Subsequently, these bodies accreted smaller, water-poor and denser pebbles over much longer timescales, when heating by Al-26 decay is less significant.
@@
Alternatively, TNOs may have formed late, only after several half-lifes of Al-26 (Bierson & Nimmo, 2019). This scenario implies that the conditions necessary for the primordial belt to form only emerged after several Myr. Subsequently, pebble accretion could proceed without significant heating constraints. However, it would then be limited by the remaining time of the gaseous disk. Formation in a ring setting, where accretion timescales are short ($\sim$1 Myr) is then the preferred scenario.
5.3 Did TNOs form by Planetesimal Accretion?
@@
An alternative to formation of TNOs by pebbles is growth among the planetesimal population itself (e.g., Kenyon & Bromley, 2004). In this scenario accretion first proceeds through a rapid runaway growth phase (Wetherill & Stewart, 1989), followed by a slower oligarchic growth phase (Kokubo & Ida, 1998). Runaway growth conditions require that the initial population of bodies of mass $m_{0}$ is dynamically quiescent (zero eccentricity), such that gravitational focusing becomes more effective for higher mass bodies. During runaway growth, a small number of bodies of mass much larger than $m_{0}$ emerge on a timescale that is a fraction of the collision timescale among the initial population, $t_{\mathrm{col,0}}$. The distribution evolves into a mass spectrum $dN/dm\propto m^{-2.5}$ (Kokubo & Ida, 1996; Barnes et al., 2009; Ormel et al., 2010b; Morishima et al., 2013). During the subsequent oligarchic growth, the most massive bodies separate from the continuous mass distribution, resulting in a flattening of the cumulative mass function. However, these oligarchs dynamically excite the planetesimal population, increasing their eccentricities and significantly slowing down growth.
@@
The similarity of the TNO mass distribution to a $p=-2.5$ power-law (see Figure 2) has been interpreted as evidence for a runaway growth origin (Ormel et al., 2010a; Kenyon & Bromley, 2012; Morishima, 2017). For the tapered-exponential distribution of Equation (2), it is natural to let $m_{0}$ coincide with the mass scale where most of the population mass resides, i.e., $m_{0}\sim 4m_{\mathrm{brk}}\approx 10^{21}\,\mathrm{g}$ (see Appendix A), corresponding to a radius of about 50  km. The initial collision timescale then evaluates
@@
where $m_{\mathrm{PB}}$ is the mass of the primordial belt and $W_{r}/r$ its fractional width. Unlike pebble accretion, planetesimal-driven growth does not need to complete before the gas disk disperses, but it must complete before the onset of the dynamical instability. The survival of the binary Jupiter Trojan Patroclus–Menoetius puts constraints on the level of dynamical interactions in the precursor primordial belt. Specifically, Nesvorný et al. (2018) find a surviving probability exponentially decaying by a factor of 10 for every 100 Myr. Studies simulating the timing of the giant planet instability are consistent with these constraints (de Sousa et al., 2020). Thus, while these constraints limit the time for collisional processing in the primordial belt, the runaway growth phase is likely to have been sufficiently rapid (Equation (22)) to establish the $p=-2.5$ mass spectrum.
However, the TNO size distribution flattens at high mass, suggesting that the high-mass bodies (Triton, Pluto, Eris) accreted significantly more mass than their low-mass counterparts. This characteristic is a hallmark of oligarchic growth, but it could also be achieved with pebble accretion (see, e.g., the lami-1p-himass model in Figure 6). Oligarchic growth, however, proceeds much slower than runaway growth and it may be challenging to meet the aforementioned timescale constraints. Alternatively, the runaway growth phase could have continued until Pluto-sized bodies, resulting in a continuous $p=-2.5$ mass spectrum, but with the subsequent stochastic implantation retaining a much higher fraction of TNO dwarf planets. In order to determine whether planetesimal or pebble accretion was the dominant formation mechanism, future simulations are needed to evaluate the scenarios under equivalent conditions. These efforts would greatly benefit from a more extensive and complete sample of TNOs.

### v2

@@ 4.1 Formation in a pebble ring
Pebble rings are ubiquitous in ALMA imagery (Andrews et al., 2018; Bae et al., 2023). The containment of particles by pressure maxima is the leading explanation for the observed particles accumulation of particles in these rings. Around the pressure maximum location $\eta\approx 0$ and the relative motion between pebbles and bigger bodies is due to turbulence. Studies have used the extent of rings in both the vertical and radial dimensions to constrain the diffusivity (the “turbulent alpha” of the gas). Specifically, this constrains the ratio of gas diffusivity over Stokes number of the particles, $\alpha_{D}/\tau_{s}$. Rosotti (2023) compiled results from a number of works that measure the radial width of continuum rings (their Table 2), and found that the inferred $\alpha_{D}/\tau_{s}$ ranged from a low of 0.04 to values exceeding unity (the latter essentially implies that particles follow the pressure distribution). However, $\alpha_{D}/\tau_{s}$ obtained from the vertical extent of the dust layer are often inferred to be smaller, e.g., $\alpha_{D}/\tau_{s}\approx 10^{-2}$ (Villenave et al., 2022). For simplicity, we adopt in the following a range in $\alpha_{D}/\tau_{s}$ between $0.01$ and $0.4$, which covers most of these measurements.
When we convert the diffusivity parameter $\alpha_{D}$ to a root mean square (rms) particle velocity, we obtain [^10] Here we have related the parameter $\alpha_{D}$, which expresses the diffusivity of the gas, to the rms velocity of the turbulent gas motions $\sigma$. The turbulent diffusivity of the gas is parameterized in terms of $\alpha$ as $D_{\mathrm{gas}}=\alpha_{D}c_{s}^{2}/\Omega$. In general, we can write $D_{\mathrm{gas}}=V_{L}^{2}t_{L}$, where $V_{L}$ and $t_{L}$ are, respectively, the velocities and turnover times of the largest eddies in the turbulent cascade. A common choice for rotating systems is $t_{L}=\Omega_{K}^{-1}$ and therefore $V_{L}=\sqrt{\alpha_{D}}c_{s}$ (Cuzzi et al., 2001). Furthermore, if the turbulence is Kolmogorov, the rms velocities of the turbulent gas is $\Delta v=\mathcal{M}c_{s}=(3/2)V_{L}$. Then, we obtain $\alpha_{D}=(2\mathcal{M}/3)^{2}=(2\sigma/3h)^{2}$.
@@
Figure 8: Constraints on the particle aerodynamic size (Stokes number $\tau_{s}$) and turbulence intensity of the gas, where $\sigma$ is identified as the root mean square turbulent gas velocity. The constraints follow from simulating the TNO size distribution (run turb-1p, black line), ALMA ring morphological analysis (blue lines; Rosotti 2023), and particle fragmentation velocity (orange lines; the adopted value for the Keplerian velocity $v_{K}$ is the circular motion at 20 au). Under the assumption that the primordial belt featured conditions similar to ALMA rings, the degeneracy between $\tau_{s}$ and $\sigma$ (or $\alpha_{D}$; Equation (16)) can be lifted and we find, $\tau_{s}\sim 0.01$, $\alpha_{D}\sim 10^{-3}$, and a fragmentation threshold of ${\sim}2\,\mathrm{m\,s}^{-1}$.
A further constraint to particle sizes in rings comes from the fragmentation barrier (Güttler et al., 2010; Birnstiel et al., 2011). When the gas is turbulent, the pebble-pebble relative velocity increases with the size (Stokes number) as $\Delta v_{\mathrm{pp}}\approx\mathcal{M}c_{s}\sqrt{2\tau_{s}}$ (Ormel & Cuzzi, 2007), [^11] This is the rms particle-particle relative velocity between two particles in the inertial range. The rms particle-planetesimal relative velocity for $\tau_{s}<1$ particles is $\Delta v=\mathcal{M}c_{s}$. where $\mathcal{M}$ is the Mach number of the turbulent flow. The aerodynamic size is then determined by the condition $\Delta v_{\mathrm{pp}}=v_{\mathrm{frag}}$, resulting in
@@
These constraints are presented in Figure 8, together with the constraint on $\tau_{s}\sigma^{3}$ that we obtained from modeling the TNO size distribution (Table 2). The constraint on the extent of ALMA rings is almost orthogonal to the constraint from the TNO size distribution modeling (this study), allowing us to put meaningful constraints on $\tau_{s}$ and $\alpha_{D}$. Together, they point to an aerodynamic size in the range of $\tau_{s}\sim[3\times 10^{-3}{-}4\times 10^{-2}]$, a turbulent diffusivity of $\alpha_{D}\sim[4\times 10^{-4}{-}2\times 10^{-3}]$, and a fragmentation threshold of $v_{\mathrm{frag}}\sim[1.7{-}3.2\,\mathrm{m\,s}^{-1}]$.
These values are similar to what was found by Jiang et al. (2024) in a recent analysis, who assumed a fragmentation threshold of $v_{\mathrm{frag}}\sim 1\,\mathrm{m\,s}^{-1}$. Our results independently suggest that the particle fragmentation threshold amount is ${\sim}2\,\mathrm{m\,s}^{-1}$ in order for particles to grow to the inferred aerodynamic sizes. A higher fragmentation threshold is possible if the growth of particles is limited by, for example, bouncing (Zsom et al., 2010). A fragmentation velocity increase by a factor of 2 would increase the derived value of the fragmentation $\alpha$ in Jiang et al. (2024) by a factor of 4 to ${\sim}4\times 10^{-4}$, in line with our findings.
@@ 4.2 A laminar disk
If pebbles are not entrained by a pressure maximum and freely drift inward, the relative velocity between planetesimals and pebbles is given by the global disk pressure gradient $\eta$. From Equation (8), $\sigma=\eta\approx 5\times 10^{-3}$ for a power-law index of pressure on radius of $2$ and from the $\tau_{s}$$\sigma$ relationship (Equation (15)), we then find an aerodynamic size of a mere $\tau_{s}=7\times 10^{-4}$, which is rather small for pebbles. The radial drift timescale
@@
already rivals the disk lifetime, suggesting that particles hardly drift. In addition, the growth of planetesimals by such small particles would be excruciatingly slow, $t_{\mathrm{gr}}\sim t_{0}/\tau_{s}\sim 50\,\mathrm{Myr}$ for the parameters listed in Equation (5), far exceeding the lifetime of the gas disk. A growth timescale ${\lesssim}10^{6}$ yr would therefore imply a (midplane) dust-to-gas ratio $Z_{\mathrm{peb}}\sim 1$, which seems unlikely for small $\tau_{s}$ particles. In summary, it is unlikely that TNOs could accrete pebbles from an inward-drifting pebble stream in the standard, laminar disk model.
In addition, pebble accretion in a laminar disk is lossy (Ormel, 2017; Lin et al., 2018). In the scenario where the birth ring is supported by pressure, all pebbles can (eventually) be accreted. That is, the efficiency is near 100%. This contrasts with the smooth disk scenario, where pebbles that fail to be accreted are lost. Consequently, the total mass in pebbles required for accretion, $m_{\mathrm{needed}}$, exceeds the mass in pebbles actually accreted ($m_{\mathrm{peb}}$). In a laminar disk, we can estimated $m_{\mathrm{needed}}$ as
@@
All these concerns would be reduced considerably when the disk is colder. Expressed in terms of the disk aspect ratio ($h\propto T^{1/2}$), $\eta\propto h^{2}$ and therefore $\tau_{s}\propto\sigma^{-3}\propto h^{-6}$ from Equation (15). For example, when $h=0.05$ (about 35 K at 20 au; see Equation (7)), $\tau_{s}\approx 6\times 10^{-3}$, a factor of 10 increase, the drift timescale reduces by a factor of 4 compared to Equation (19), the growth timescale may fall below the lifetime of the disk, and the required pebble mass, $m_{\mathrm{needed}}$, to values below $500\,m_{\oplus}$. These numbers are within the realm of pebble accretion models (Lambrechts et al., 2019), although on the high side. A question that arises, however, is where these $500\,m_{\oplus}$ in solids would end up as the Solar System’s giant planets, while enriched, are not dominated in metals.
@@ 5.2 Formation time constraints
Recently, Cañas et al. (2024) also investigated the growth of TNOs by pebble accretion. Their approach is complementary to ours. In particular, their study emphasizes the composition and internal density of TNOs. A key constraint is that small TNOs of radius $50{-}100$ km could not have formed too quickly as radiogenic heating (primarily Al-26 decay at a half-life of 0.7 Myr Norris et al. 1983) would have melted these bodies and reduced their porosity. This outcome is in contrast with observationally inferred densities of TNOS, implying that these TNOs have retained significant porosity (Bierson & Nimmo, 2019).
@@
Cañas et al. (2024) addressed this conundrum by hypothesizing that the initial distribution of TNOs formed through streaming instabilities dominated by large, ice-rich pebbles. Processes such as photodesorption could have resulted in a compositional dichotomy where ice is effectively transferred to the larger particles that settled to the midplane layers that were shielded from stellar UV irradiation. Cañas et al. (2024) argue that streaming instability would favor the larger ice-rich pebbles over the smaller Al-26 rich ones. Hence, the first-generation planetesimals were low in Al-26 and the early heating problem outlined by Bierson & Nimmo (2019) is avoided. Subsequently, these bodies accreted smaller, more water-poor, and denser pebbles over much longer timescales when heating by Al-26 decay is less significant.
@@
Alternatively, TNOs may have formed late, only after several half-lives of Al-26 (Bierson & Nimmo, 2019). This scenario implies that the conditions necessary for the primordial belt to form only emerged after several million years. Subsequently, pebble accretion could proceed without significant heating constraints. However, it would then be limited by the remaining time of the gaseous disk. Formation in a ring setting, where accretion timescales are short ($\sim$1 Myr) is then the preferred scenario.
5.3 TNOs formation by planetesimal accretion
@@
An alternative to formation of TNOs by pebbles is growth among the planetesimal population itself (e.g., Kenyon & Bromley, 2004). In this scenario accretion first proceeds through a rapid runaway growth phase (Wetherill & Stewart, 1989), followed by a slower oligarchic growth phase (Kokubo & Ida, 1998). Runaway growth conditions require that the initial population of bodies of mass $m_{0}$ is dynamically quiescent (zero eccentricity), such that gravitational focusing becomes more effective for higher mass bodies. During runaway growth, a small number of bodies of mass much higher than $m_{0}$ emerge on a timescale that is a fraction of the collision timescale among the initial population, $t_{\mathrm{col,0}}$. The distribution evolves into a mass spectrum $dN/dm\propto m^{-2.5}$ (Kokubo & Ida, 1996; Barnes et al., 2009; Ormel et al., 2010b; Morishima et al., 2013). During the subsequent oligarchic growth, the most massive bodies separate from the continuous mass distribution, resulting in a flattening of the cumulative mass function. However, these oligarchs dynamically excite the planetesimal population, increasing their eccentricities and significantly slowing down growth.
@@
The similarity of the TNO mass distribution to a $p=-2.5$ power law (see Figure 2) has been interpreted as evidence for a runaway growth origin (Ormel et al., 2010a; Kenyon & Bromley, 2012; Morishima, 2017). For the tapered-exponential distribution of Equation (2), it is natural to let $m_{0}$ coincide with the mass scale where most of the population mass resides (i.e., $m_{0}\sim 4m_{\mathrm{brk}}\approx 10^{21}\,\mathrm{g}$; see Appendix A), corresponding to a radius of about 50  km. The initial collision timescale then evaluates
@@
where $m_{\mathrm{PB}}$ is the mass of the primordial belt and $W_{r}/r$ its fractional width. Unlike pebble accretion, planetesimal-driven growth does not need to be completed before the gas disk disperses, but it must conclude before the onset of the dynamical instability. The survival of the binary Jupiter trojan Patroclus–Menoetius puts constraints on the level of dynamical interactions in the precursor primordial belt. Specifically, Nesvorný et al. (2018) find a surviving probability exponentially decaying by a factor of 10 for every 100 Myr. Studies simulating the timing of the giant planet instability are consistent with these constraints (de Sousa et al., 2020). Thus, while these constraints limit the time for collisional processing in the primordial belt, the runaway growth phase is likely to have been sufficiently rapid (Equation (22)) to establish the $p=-2.5$ mass spectrum.
However, the TNO size distribution flattens at high mass, suggesting that the high-mass bodies (Triton, Pluto, Eris) accreted significantly more mass than their low-mass counterparts. This characteristic is a hallmark of oligarchic growth, but it could also be achieved with pebble accretion (see, e.g., the lami-1p-himass model in Figure 6). Oligarchic growth, however, proceeds much more slowly than runaway growth and it may be challenging to meet the aforementioned timescale constraints. Alternatively, the runaway growth phase could have continued until Pluto-sized bodies formed, resulting in a continuous $p=-2.5$ mass spectrum, but with the subsequent stochastic implantation retaining a much higher fraction of TNO dwarf planets. In order to determine whether planetesimal or pebble accretion was the dominant formation mechanism, future simulations are needed to evaluate the scenarios under equivalent conditions. These efforts would greatly benefit from a more extensive and complete sample of TNOs.